A jet of water of 10 mm diameter impinges normally on a smooth flat
plate of 50 N in weight as depicted below. The water is deflected and
flows out along the plate and the plate remains stationary. Assume the
friction loss and head difference are negligible, determine the flow rate.
1
Expert's answer
2020-04-27T10:33:35-0400

Let the flow rate be "v \frac{m}{sec}"

Now

Rate of change of momentum = weight = 50 N


"So\\\frac{dm}{dt}\times v= 50"

"(\rho\times\pi \ r^2\times v) \times v= 50\\1000\times 3.14\times(5\times10^{-3})^2\times v^2=50\\v^2= 636.94\\v=25.24\ \frac{m}{sec}"